# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ 1 Introduction
Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023) addresses the scarcity issue by providing a framework for valid statistical analysis using predictions from black-box ML models. By combining ML-predicted and gold-standard outcomes, PPI and its variants (Angelopoulos et al., 2024; Zrnic and Candès, 2024; Zrnic, 2025) use the abundance of predictions to reduce variance while relying on the accuracy of labeled [^1] Throughout the paper, we will use the terms “labeled” and “gold-standard” interchangeably. data to control bias.
In this work, we adapt PPI to the estimation of multiple outcome means in compound estimation settings. Many applications of PPI naturally involve parallel statistical problems that can be solved simultaneously. For instance, several PPI methods (Angelopoulos et al., 2024; Fisch et al., 2024) have shown improvements in estimating the fraction of spiral galaxies using predictions on images from the Galaxy Zoo 2 dataset (Willett et al., 2013). While these methods focus on estimating a single overall fraction, a richer analysis emerges from partitioning galaxies based on metadata (such as celestial coordinates or pre-defined bins) and estimating the fraction of galaxies within each partition. This compound estimation approach enables more granular scientific inquiries that account for heterogeneity across galaxy clusters and spatial locations (Nair and Abraham, 2010).
We demonstrate, both theoretically and empirically, the benefits of solving multiple mean estimation problems simultaneously. Our approach builds on the empirical Bayes principle of sharing information across problems (Robbins, 1956; Efron, 2010) as exemplified by James-Stein shrinkage (James and Stein, 1961; Xie et al., 2012). The connection between modern and classical statistical ideas allows us to perform within problem PPI estimation in the first place, followed by a shrinkage process reusing the ML predictions in an adaptive manner, which becomes possible through borrowing information across problems. Our contributions are as follows:
@@
    We conduct extensive experiments on both synthetic and real-world datasets. Our experiments demonstrate PAS’s applicability to large-scale problems with deep learning models, showing improved estimation accuracy compared to other estimators (e.g., classical, PPI++).
@@ 2.1 Prediction-Powered Inference (PPI)
The PPI framework considers a setting where we have access to a small number of labeled data points $(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a large number of unlabeled covariates $(\tilde{X}_{i})_{i=1}^{N}\in(\mathcal{X})^{N}$, where $\mathcal{X}$ and $\mathcal{Y}$ represent the covariate and outcome space, respectively. The data points are drawn iid from a joint distribution $\mathbb{P}_{XY}$. [^2] To be concrete: $(X_{i},Y_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$ and $(\tilde{X}_{i},\tilde{Y}_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$, but $\tilde{Y}_{i}$ is unobserved. We are also given a black-box predictive model $f:\mathcal{X}\to\mathcal{Y}$ that is independent of the datasets (e.g., pre-trained on similar but unseen data). For mean estimation with $\mathcal{Y}\subset\mathbb{R}$, the goal is to leverage the predicted outcomes $f(X_{i})$ to improve the estimation of $\theta:=\mathbb{E}[Y_{i}]$. Some simple estimators take the form of the following aggregated (summary) statistics
$\bar{Y}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\>\,$ ${\color[rgb]{.75,.75,.75}\tilde{Y}:=\frac{1}{N}\sum_{i=1}^{N}\tilde{Y}_{i},}$ (1)
@@
Above, $\bar{Y}$ is the classical estimator, [^3] From now on, we will use the term “classical estimator” to refer to the sample average of the labeled outcomes. $\bar{Z}^{f},\tilde{Z}^{f}$ are the prediction means on the labeled and unlabeled data, and $\tilde{Y}$ (greyed out) is unobserved. The vanilla PPI estimator is defined as,
@@
Both definitions represent $\hat{\theta}^{\text{PPI}}$ in the form of a baseline estimator plus a correction term. In the first representation, the baseline estimator is the unbiased classical estimator $\bar{Y}$, while the correction term has expectation 00 and attempts to reduce the variance of $\bar{Y}$. In the second representation, the baseline estimator is the prediction mean on unlabeled data $\tilde{Z}^{f}$ (which in general may be biased for $\theta$), while the correction term removes the bias of $\tilde{Z}^{f}$ by estimating the bias of the ML model $f$ on the labeled dataset. Writing $\hat{\theta}^{\text{PPI}}=\tfrac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i})+\tfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))$, we find
$\mathbb{E}[\hat{\theta}^{\text{PPI}}]$ $=\mathbb{E}[Y_{i}]=\theta,$
@@
The classical estimator $\bar{Y}$ which only uses labeled data is unbiased for $\theta$ and has variance (and MSE) equal to $n^{-1}\mathrm{Var}[Y_{i}]$, while the MSE of $\hat{\theta}^{\text{PPI}}$ in (3) may be arbitrarily large when the predictive model is inaccurate.
@@
which recovers the classical estimator when $\lambda=0$ and the vanilla PPI estimator when $\lambda=1$. For all values of $\lambda$, $\,\hat{\theta}^{\text{PPI}}_{\lambda}$ is unbiased, so if we select the $\lambda$ that minimizes $\mathrm{Var}[\hat{\theta}^{\text{PPI}}_{\lambda}]$, we can improve our estimator over both the classical estimator and vanilla PPI. Such an estimator is defined as the Power-Tuned PPI (PT [^4] We will use the term “PPI++” for the larger framework, while “PT” refers to the specific estimator.) estimator $\hat{\theta}^{\mathrm{PT}}:=\hat{\theta}^{\text{PPI}}_{\lambda^{*}}$, where
$\lambda^{*}:=\operatorname*{arg\,min}_{\lambda\in[0,1]}\mathrm{Var}[\hat{\theta}^{\text{PPI}}_{\lambda}].$
Power-tuning will be one of the building blocks of PAS in Section 4.
@@
Consider the ideal scenario for PPI with $N=\infty$ (that is, the unlabeled dataset is much larger than the labeled dataset) so that $\tilde{Z}^{f}\equiv\mathbb{E}[f(\tilde{X}_{i})]$. Even then, the MSE of $\hat{\theta}^{\text{PPI}}$ in (3) is always lower bounded [^5] The same lower bound also applies to power-tuned PPI $\hat{\theta}^{\mathrm{PT}}$. by $\mathbb{E}[\mathrm{Var}[Y_{i}\mid X_{i}]]/n$ and the lower bound is attained for the perfect ML predictor $f(\cdot)\equiv\mathbb{E}\!\left[Y_{i}\mid X_{i}=\cdot\right]$. In words, if $Y_{i}$ is not perfectly predictable from $X_{i}$, then PPI applied to a labeled dataset of fixed size $n$ must have nonnegligible MSE. By contrast, for $N=\infty$, the prediction mean of unlabeled data $\tilde{Z}^{f}$ has zero variance and MSE equal to the squared bias $(\mathbb{E}[f(X_{i})]-\theta_{i})^{2}$. Thus if the predictor satisfies a calibration-type property that $\mathbb{E}[f(X_{i})]\approx\mathbb{E}[Y_{i}]$ (which is implied by, but much weaker than the requirement $f(X_{i})\approx Y_{i}$), then the MSE of $\tilde{Z}^{f}$ could be nearly 00. By contrast, PPI (and PPI++) can only partially capitalize on such a predictor $f(\cdot)$.
@@ 2.2 The Compound Mean Estimation Setting
In this section, we introduce the problem setting that PAS is designed to address—estimating the mean of $m>1$ parallel problems with a single black-box predictive model $f$ [^6] Our proposal also accommodates using separate predictors $\{f_{j}\}_{j=1}^{m}$ for each problem. To streamline exposition, we focus on the practical scenario where a single (large) model (e.g., an LLM or vision model) can handle multiple tasks simultaneously (Radford et al., 2019; He et al., 2022). For the $j$-th problem, where $j\in[m]:=\{1,\ldots,m\}$, we observe a labeled dataset $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ observations and an unlabeled dataset $(\tilde{X}_{ij})_{i=1}^{N_{j}}$ with $N_{j}\in\mathbb{N}$ observations. We start with modeling heterogeneity across problems.
@@
For each problem $j\in[m]$, we assume that the joint distribution of $(f(X_{ij}),Y_{ij})$ has finite second moments conditional on $\eta_{j}$ and that
$\begin{bmatrix}f(X_{ij})\\ Y_{ij}\end{bmatrix}\,\big{|}\,\eta_{j}\,\,{\buildrel\text{iid}\over{\sim}\,}\,\mathbb{F}_{j}\left(\begin{bmatrix}\mu_{j}\\ \theta_{j}\end{bmatrix},\begin{bmatrix}\tau_{j}^{2}&\rho_{j}\tau_{j}\sigma_{j}\\ \rho_{j}\tau_{j}\sigma_{j}&\sigma_{j}^{2}\end{bmatrix}\right),\;i\in[n_{j}],\,\,$ (7)
where $\mathbb{F}_{j}$, $\mu_{j},\theta_{j},\rho_{j},\sigma_{j}^{2},\tau^{2}_{j}$ are functions of $\eta_{j}$. Conditional on $\eta_{j}$, the unlabeled predictions $f(\tilde{X}_{i^{\prime}j})$, $i^{\prime}\in[N_{j}]$, are also iid, independent of the labeled dataset and identically distributed with $f(X_{ij})$. In the notation of (7), $\mathbb{F}_{j}$ represents a unspecified distribution satisfying the moment constraints in (6) and
@@
Figure 1: We instantiate the model described in Example 2.3 with $m=10$ problems, each has $n_{j}=10$ labeled and $N_{j}=20$ unlabeled data (in different colors). (Top) Labeled data $(X_{ij},Y_{ij})_{j=1}^{n_{j}}$ with the classical estimator $\bar{Y}_{j}$ shown for each problem. (Bottom) We apply a flawed predictor $f(x)=|x|$ to the unlabeled covariates and visualize $(X_{ij},f(X_{ij}))_{j=1}^{N_{j}}$ as well as the prediction mean $\tilde{Z}_{j}^{f}$.
Similar to Eq. (1), we define the aggregated statistics $\bar{Y}_{j},\bar{Z}_{j}^{f},\tilde{Z}^{f}_{j}$ for each $j\in[m]$. Following prior work, [^7] For EB, examples include Xie et al. (2012); Soloff et al. (2024); for PPI, see recent works like Fisch et al. (2024). we treat $\tau_{j}$, $\sigma_{j}$, $\rho_{j}$ as known in our development. The rationale is that estimation of $\tau_{j},\sigma_{j},\rho_{j}$ is a second-order concern in our setup of mean estimation and treating these as known facilitates exposition. In practice, such as in our numerical experiments with real-world datasets, we use sample-based estimates of $\tau_{j},\sigma_{j},$ and $\rho_{j}$ (see Section D.1).
@@
For each problem $j$, let $\eta_{j}\sim\mathcal{U}[-1,1]$. We think of $\eta_{j}$ as both indexing the problems, and generating heterogeneity across problems. The $j$-th dataset is generated via (with constants $c=0.05,\psi=0.1$),
@@
In Figure 1, we visualize realizations from this model with $m=10$ problems, $n_{j}=10$ labeled observations, and $N_{j}=20$ unlabeled observations for each problem. We apply a flawed predictor $f(x)=|x|$. The classical estimator $\bar{Y}_{j}$ and the prediction mean $\tilde{Z}^{f}_{j}$ deviate from each other. Nevertheless, $\tilde{Z}^{f}_{j}$ contains information that can help us improve upon $\bar{Y}_{j}$ as an estimator of $\theta_{j}$ by learning from within problem (PPI, PPI++, this work) and across problem (this work) structure. We emphasize that as specified in (7), our approach only requires modeling the first and second moments of the joint distribution of $(f(X_{ij}),Y_{ij})$. For instance, in this synthetic model, $\theta_{j}=\eta_{j}^{2}$ and $\sigma_{j}^{2}=4\eta_{j}^{2}\psi^{2}+c$, while $\mu_{j}$, $\tau_{j}^{2}$ and $\gamma_{j}$ also admit closed-form expressions in terms of $\eta_{j}$ when the predictor takes the form $f(x)=|x|$ or $f(x)=x^{2}$ (see Section D.2).
To conclude this section, we define the compound risk (Robbins, 1951; Jiang and Zhang, 2009) for any estimator $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{m})^{\intercal}$ as the expected squared error loss averaged over problems,
@@
$\text{where}\quad\ell_{m}(\boldsymbol{\hat{\theta}},\boldsymbol{\theta})$ $:=\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}-\theta_{j})^{2}.$ (10)
@@
3 Statistical Guiding Principles and Connections to Prior Work
In this section, we illustrate both the statistical guiding principles of our approach and some connections to prior work through the following stylized Gaussian model:
Sampling: $\hat{\theta}^{\text{cl}}=\theta+(\xi+\varepsilon),$ $\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2}),\quad\quad\,\,\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}).$
Prior: $\theta\sim\mathcal{N}(0,\sigma_{\theta}^{2}),$ $\phi\sim\mathcal{N}(0,\sigma_{\phi}^{2}),\quad\quad\mathrm{Corr}\!\left[\theta,\phi\right]=\rho.$

### v2

@@ 1 Introduction
Prediction-Powered Inference (PPI) (Angelopoulos et al., 2023) addresses the scarcity issue by providing a framework for valid statistical analysis using predictions from black-box ML models. By combining ML-predicted and gold-standard outcomes, PPI and its variants (Angelopoulos et al., 2024; Zrnic & Candès, 2024; Zrnic, 2025) use the abundance of predictions to reduce variance while relying on the accuracy of labeled [^1] Throughout the paper, we use the terms “labeled” and “gold-standard” interchangeably. data to control bias.
In this work, we adapt PPI to the estimation of multiple outcome means in compound estimation settings. Many applications of PPI naturally involve parallel statistical problems that can be solved simultaneously. For instance, several PPI methods (Angelopoulos et al., 2024; Fisch et al., 2024) have shown improvements in estimating the fraction of spiral galaxies using predictions on images from the Galaxy Zoo 2 dataset (Willett et al., 2013). While these methods focus on estimating a single overall fraction, a richer analysis emerges from partitioning galaxies based on metadata (such as celestial coordinates or pre-defined bins) and estimating the fraction of galaxies within each partition. This compound estimation approach enables more granular scientific inquiries that account for heterogeneity across galaxy clusters and spatial locations (Nair & Abraham, 2010).
We demonstrate, both theoretically and empirically, the benefits of solving multiple mean estimation problems simultaneously. Our approach builds on the empirical Bayes (EB) principle of sharing information across problems (Robbins, 1956; Efron, 2010) as exemplified by James-Stein shrinkage (James & Stein, 1961; Xie et al., 2012). The connection between modern and classical statistical ideas allows us to perform within problem PPI estimation in the first place, followed by a shrinkage step reusing the ML predictions in an adaptive way, which becomes possible through borrowing information across problems. Our contributions are:
@@
    We conduct extensive experiments on both synthetic and real-world datasets. Our experiments demonstrate PAS’s applicability to large-scale problems with deep learning models, showing improved estimation accuracy compared to other classical and modern baselines.
@@ 2.1 Prediction-Powered Inference (PPI)
The PPI framework considers a setting where we have access to a small number of labeled data points $(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a large number of unlabeled covariates $(\tilde{X}_{i})_{i=1}^{N}\in(\mathcal{X})^{N}$, where $\mathcal{X}$ and $\mathcal{Y}$ represent the covariate and outcome space, respectively. The data points are drawn iid from a joint distribution $\mathbb{P}_{XY}$. [^2] To be concrete: $(X_{i},Y_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$ and $(\tilde{X}_{i},\tilde{Y}_{i})\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}_{XY}$ independently, but $\tilde{Y}_{i}$ is unobserved. We are also given a black-box predictive model $f:\mathcal{X}\to\mathcal{Y}$ that is independent of the datasets (e.g., pre-trained on similar but unseen data). For mean estimation with $\mathcal{Y}\subset\mathbb{R}$, the goal is to leverage the predicted outcomes $f(X_{i})$ to improve the estimation of $\theta:=\mathbb{E}[Y_{i}]$. Some simple estimators take the form of the following aggregated (summary) statistics
$\bar{Y}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i},\>\,$ ${\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}\tilde{Y}:=\frac{1}{N}\sum_{i=1}^{N}\tilde{Y}_{i},}$ (1)
@@
Above, $\bar{Y}$ is the classical estimator, [^3] From now on, we will use the term “classical estimator” to refer to the sample average of the labeled outcomes. $\bar{Z}^{f},\tilde{Z}^{f}$ are the prediction means on the labeled and unlabeled data, and $\tilde{Y}$ (grayed out) is unobserved. The vanilla PPI estimator is defined as,
@@
Both definitions represent $\hat{\theta}^{\text{PPI}}$ in the form of a baseline estimator plus a correction term. In the first representation, the baseline estimator is the unbiased classical estimator $\bar{Y}$, while the correction term has expectation 00 and attempts to reduce the variance of $\bar{Y}$. In the second representation, the baseline estimator is the prediction mean on unlabeled data $\tilde{Z}^{f}$ (which in general may be biased for $\theta$), while the correction term removes the bias of $\tilde{Z}^{f}$ by estimating the bias of the ML model $f$ on the labeled dataset. Writing $\hat{\theta}^{\text{PPI}}=\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i})+\tfrac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))$, we find that $\mathbb{E}[\hat{\theta}^{\text{PPI}}]=\mathbb{E}[Y_{i}]=\theta$ and
@@
The classical estimator $\bar{Y}$ which only uses labeled data is unbiased for $\theta$ and has variance (and MSE) equal to $\tfrac{1}{n}\mathrm{Var}[Y_{i}]$.
@@
which recovers the classical estimator when $\lambda=0$ and the vanilla PPI estimator when $\lambda=1$. For all values of $\lambda$, $\,\hat{\theta}^{\text{PPI}}_{\lambda}$ is unbiased, so if we select the $\lambda$ that minimizes $\mathrm{Var}[\hat{\theta}^{\text{PPI}}_{\lambda}]$, we can improve our estimator over both the classical estimator and vanilla PPI. Such an estimator is defined as the Power-Tuned (PT) PPI [^4] We use the term “PPI++” for the broader framework, while “PT” refers to the specific estimator. estimator $\hat{\theta}^{\mathrm{PT}}:=\hat{\theta}^{\text{PPI}}_{\lambda^{*}}$, where we pick $\lambda^{*}$ that minimizes the variance (and thus the MSE) of $\hat{\theta}^{\text{PPI}}_{\lambda}$. We will revisit PT as one of the building blocks of our proposed PAS estimator in Section 4.
@@
Consider the ideal scenario for PPI with $N=\infty$ (that is, the unlabeled dataset is much larger than the labeled dataset) so that $\tilde{Z}^{f}\equiv\mathbb{E}[f(\tilde{X}_{i})]$. Even then, the MSE of $\hat{\theta}^{\text{PPI}}$ in (3) is always lower bounded [^5] The same lower bound also applies to power-tuned PPI $\hat{\theta}^{\text{PT}}$. by $\tfrac{1}{n}\mathbb{E}[\mathrm{Var}[Y_{i}\mid X_{i}]]$ and the lower bound is attained by the perfect ML predictor $f(\cdot)\equiv\mathbb{E}\!\left[Y_{i}\mid X_{i}=\cdot\right]$. In words, if $Y_{i}$ is not perfectly predictable from $X_{i}$, then PPI applied to a labeled dataset of fixed size $n$ must have non-negligible MSE. By contrast, for $N=\infty$, the prediction mean of unlabeled data $\tilde{Z}^{f}$ has zero variance and MSE equal to the squared bias $(\mathbb{E}[f(X_{i})]-\theta_{i})^{2}$. Thus if the predictor satisfies a calibration-type property that $\mathbb{E}[f(X_{i})]\approx\mathbb{E}[Y_{i}]$ (which is implied by, but much weaker than the requirement $f(X_{i})\approx Y_{i}$), then the MSE of $\tilde{Z}^{f}$ could be nearly 00. By contrast, PPI (and PPI++) can only partially capitalize on such a predictor $f(\cdot)$.
@@ 2.2 The Compound Mean Estimation Setting
In this section, we introduce the problem setting that PAS is designed to address—estimating the mean of $m>1$ parallel problems with a single black-box predictive model $f$ [^6] Our proposal also accommodates using separate predictors $\{f_{j}\}_{j=1}^{m}$ for each problem. To streamline exposition, we focus on the practical scenario where a single (large) model (e.g., a large language or vision model) can handle multiple tasks simultaneously (Radford et al., 2019; He et al., 2022). For the $j$-th problem, where $j\in[m]:=\{1,\ldots,m\}$, we observe a labeled dataset $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$ with $n_{j}\in\mathbb{N}$ samples and an unlabeled dataset $(\tilde{X}_{ij})_{i=1}^{N_{j}}$ with $N_{j}\in\mathbb{N}$ samples. We start with modeling heterogeneity across problems.
@@
For each problem $j\in[m]$, we assume that the joint distribution of $(f(X_{ij}),Y_{ij})$ has finite second moments conditional on $\eta_{j}$ and for $i\in[n_{j}]$
$\begin{bmatrix}f(X_{ij})\\ Y_{ij}\end{bmatrix}\,\big{|}\,\eta_{j}\,\,{\buildrel\text{iid}\over{\sim}\,}\,\mathbb{F}_{j}\left(\begin{bmatrix}\mu_{j}\\ \theta_{j}\end{bmatrix},\begin{bmatrix}\tau_{j}^{2}&\rho_{j}\tau_{j}\sigma_{j}\\ \rho_{j}\tau_{j}\sigma_{j}&\sigma_{j}^{2}\end{bmatrix}\right),$ (7)
where $\mathbb{F}_{j}$, $\mu_{j},\theta_{j},\rho_{j},\sigma_{j}^{2},\tau^{2}_{j}$ are functions of $\eta_{j}$. Conditional on $\eta_{j}$, the unlabeled predictions $f(\tilde{X}_{i^{\prime}j})$, $i^{\prime}\in[N_{j}]$, are also iid, independent of the labeled dataset and identically distributed with $f(X_{ij})$. In the notation of (7), $\mathbb{F}_{j}$ represents an unspecified distribution satisfying the moment constraints in (6) and
@@
Figure 1: We instantiate the model described in Example 2.3 with $m=10$ problems, each has $n_{j}=10$ labeled and $N_{j}=20$ unlabeled data (we use different colors for all 10 problems). (Top) Labeled data $(X_{ij},Y_{ij})_{j=1}^{n_{j}}$ with the classical estimator $\bar{Y}_{j}$ shown for each problem. (Bottom) We apply a flawed predictor $f(x)=|x|$ to the unlabeled covariates and visualize $(X_{ij},f(X_{ij}))_{j=1}^{N_{j}}$ as well as the prediction mean $\tilde{Z}_{j}^{f}$.
Similar to Eq. (1), we define the aggregated statistics $\bar{Y}_{j},\bar{Z}_{j}^{f},\tilde{Z}^{f}_{j}$ for each $j\in[m]$. To facilitate exposition, following prior work, [^7] For EB, examples include Xie et al. (2012); Soloff et al. (2024); for PPI, see recent works like Fisch et al. (2024). we treat the second moments $\tau_{j}^{2}$, $\sigma_{j}^{2}$, $\gamma_{j}$ as known until the end of Section 5. In Section 6, we extend our methods to allow for unknown $\tau_{j}^{2}$, $\sigma_{j}^{2}$, and $\gamma_{j}$.
@@
For each problem $j$, let $\eta_{j}\sim\mathcal{U}[-1,1]$. We think of $\eta_{j}$ as both indexing the problems and generating heterogeneity across problems. The $j$-th dataset is generated via (with constants set to $c=0.05,\psi=0.1$),
@@
In Figure 1, we visualize realizations from this model with $m=10$ problems, $n_{j}=10$ labeled observations, and $N_{j}=20$ unlabeled observations for each problem. We apply a flawed predictor $f(x)=|x|$. The classical estimator $\bar{Y}_{j}$ and the prediction mean $\tilde{Z}^{f}_{j}$ deviate from each other. Nevertheless, $\tilde{Z}^{f}_{j}$ contains information that can help us improve upon $\bar{Y}_{j}$ as an estimator of $\theta_{j}$ by learning from within problem (PPI, PPI++, this work) and across problem (this work) structure. We emphasize that, as specified in (7), our approach only requires modeling the first and second moments of the joint distribution of $(f(X_{ij}),Y_{ij})$. For instance, in this synthetic model, $\theta_{j}=\eta_{j}^{2}$ and $\sigma_{j}^{2}=4\eta_{j}^{2}\psi^{2}+c$, while $\mu_{j}$, $\tau_{j}^{2}$ and $\gamma_{j}$ also admit closed-form expressions in terms of $\eta_{j}$ when the predictor takes the form $f(x)=|x|$ or $f(x)=x^{2}$ (see Section E.1).
To conclude this section, we define the compound risk (Robbins, 1951; Jiang & Zhang, 2009) for any estimator $\boldsymbol{\hat{\theta}}=(\hat{\theta}_{1},\ldots,\hat{\theta}_{m})^{\intercal}$ as the expected squared error loss averaged over problems,
@@
$\text{with}\quad\ell_{m}(\boldsymbol{\hat{\theta}},\boldsymbol{\theta})$ $:=\frac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}-\theta_{j})^{2}.$ (10)
@@
3 Statistical Guiding Principles & Prior Work
In this section, we illustrate both the statistical guiding principles of our approach and some connections to prior work [^8] We provide further connections in Section A.1. through the following stylized Gaussian model:
Sampling: $\!\!\!\!\hat{\theta}^{\text{cl}}=\theta+(\xi+\varepsilon),\,\xi\sim\mathcal{N}(0,\sigma_{\xi}^{2}),\,\varepsilon\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2}).$
Prior: $\!\!\!\!\theta\sim\mathcal{N}(0,\sigma_{\theta}^{2}),\,\phi\sim\mathcal{N}(0,\sigma_{\phi}^{2}),\,\mathrm{Corr}\!\left[\theta,\phi\right]=\rho.$